Use Gauss-Jordan row reduction to solve the given system of equations. HINT [See Examples 1-6.] (If there is no solution, enter NO SOLUTION. If the system is dependent, express your answer in terms of z, where

x = x(z)

and

y = y(z).)

3x y z = 0
x + y + z = 4
(x, y, z) =
Answer #1

On adding the given two equations, we will get:

4x = 4

x = 1

On putting x = 1 in the first equation:

3 - y - z = 0

y + z = 3

y = 3 - z

Hence the solution will be:

(x, y, z) = (1, 3 - z, z)